SOLVE AN PLEASE SHOW WORK
30x^2 - 9x - 3=0
Any help would be greatly appreciated. Thank you for your kindness!

Answer by checkley71(8403)

30X~2-9X-3=0 OR (5X+1)(6X-3)=0 OR 5X=-1 OR X=-1/5 & 6X=3 OR X=1/2
PROOF 30(-1/5)~2-9(-1/5)-3=0 OR 30/25+9/5-3=0 OR 6/5+9/5-3=0 OR 15/5-3=0 OR
3-3=0 OR 0=0